[Brahms-l] RHIC schedule

From: flemming videbaek <videbaek@rcf.rhic.bnl.gov>
Date: Mon Dec 13 2004 - 14:10:45 EST
Dear Collaborator,
I just got the most recent update from Kin Yip - his message is first,
followed by an extract from his pointed news list.

As you can see the overall delay is considerable ~ 3 weeks. Your guess is probably as good as mine,
the firsts colliding beams seems not to happen until after new year.

flemming



====== Acting Scheduling Physicist -YIP (ie. A-SPY) news channel ========

(0) There will be NO Time Meeting tomorrow (Dec. 14, 2004).  

Please don't feel too disappointed :-)

***********************************************************************************************

(1) Brief schedule :

I have made up the first draft of the schedule:
http://www.c-ad.bnl.gov/kinyip/SchedPhys/RHIC_shorts_2004/

(if you haven't checked before, it's at the top of this webpage inside a "box".)

I'll write a couple lines every now and then this week in the above "box" if there
is something worth knowing about.

If all go well/expected (including the Blue ring being available for cool-down first),  
the Blue ring could be colded on ~Dec. 22 and the Yellow ring could be colded 
4/5 days later.  The magnet testing and then RHIC start-up would follow. 
Let's hope that the Yellow ring has no more problem.

***********************************************************************************************

(2) Brief description of the "short" problem in 12 o'clock near Q3 
(mostly from George):

The problem for the Yellow ring ~12 o'clock was a bus to bus short that
started an arc when the shunt busses had the highest voltage between
them( appox. 200 Volts ). This arc then caused the conductors to melt
and cause the busses to have a high resistance. 

***********************************************************************************************

(3) Time lost:

So, this shutdown is more or less like 3 weeks (instead of the original 2 weeks)
if you count from Dec. 3, 2004 (Friday).

Fortunately (or not too unfortunately), the power and thus financial consumption in 
this incomplete shutdown mode is about half of what it is during normal RHIC running.  
So, we kind of lost ~1.5 weeks financially.





--------------------------------------------------------------------------------

Dec. 13, 2004:

1st estimated schedule looks like the following:

  a.. Tue night (Dec. 14): George & Co. finish the electrical work; 
  b.. Wed./Thur.: Gary & Co. hand over the process line to the Cryogenic Group; 
  c.. Sat./Sun.: All the vessels (that we have opened) should be closed and ready for vacuum pumping down; 
  d.. Sun. (Dec. 19): Blue ring ready for the Cryogenic System to cool down (which may take 3-4 days); 
  e.. Next Tue. (Dec. 21): Yellow ring ready for the Cryogenic System to cool down (which may take 5-6 days); 
  f.. Priority is given to the Yellow ring.  But if the Blue ring is available, we cool the Blue immediately.  The  earliest day for the colded Blue ring is ~ Dec. 22 and the earliest day for the colded Yellow ring is ~Dec. 26. 
  g.. These would be immediately followed by magnet testing (hopefully the Yellow ring has no more problem !) and RHIC start-up. 

--------------------------------------------------------------------------------

Evening  Dec. 12, 2004:

George thinks that he should be done with his electrical work by Tuesday night. I'm gathering information to make the 1st estimated schedule for restarting the FY05 RHIC run. 


--------------------------------------------------------------------------------

Evening  Dec. 11, 2004:

12 o'clock area: George & Co. have found the short in the Q3 splice.  Preparation is being done and repair will start tomorrow (Sunday !).

For 10 o'clock, we'll be closing up the 4 interconnects that we have opened.  It is said that welding to close those Helium Bellows may start next Tue. or Wed. 


--------------------------------------------------------------------------------

Evening  Dec. 10, 2004:

All 4 inter-connect vessels (between D6 and D8) at 10 o'clock have been cut open.   No obvious damage which might have led to short has been found.  But it's believed that we might be able to work around it even if we don't find the origin of the short at 10 o'clock.

The Blue ring at 12 o'clock has been warmed up.   The "cutting" may be done tonight and George and his people would start to deal with the 12 o'clock problem (short leading to circuit being opened) then.


--------------------------------------------------------------------------------

Morning  Dec. 10, 2004:

No obvious spot for short was found last night.   A 3rd vessel is being cut open.   At ~2 pm, they are going to do a high-pot to try to recreate the short.   Apparently, the "short" disappeared after warming up.


--------------------------------------------------------------------------------

~17:00  Dec. 9, 2004:

No obvious "short" is found even after the 2nd vessel is cut open.  So, George is diagnosing the problem "electrically" tonight.

----------------------------------------------------------------
Flemming Videbaek
Physics Department
Brookhaven National Laboratory

e-mail: videbaek@bnl.gov
phone: 631-344-4106



_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Dec 13 14:09:05 2004

This archive was generated by hypermail 2.1.8 : Mon Dec 13 2004 - 14:09:29 EST